For part 3 of 3, just enter magnitude (positve number) A 1720 kg car has a speed of 10.3 m/s when it hits a tree. The tree doesn’t move and the car comes to rest. Find the change in kinetic energy of the car. Answer in units of J. 008 (part 2 of 3) 10.0 points Find the amount of work done by the car as its front is pushed in. Answer in units of J. 009 (part 3 of 3) 10.0 points Find the magnitude of the force that pushed the front of the car in by 50 cm . Answer in units of N.


1
Expert's answer
2020-12-10T11:10:08-0500

mass of the car, "m = 1720 kg"

initial speed, "u = 10.3 m\/s"

When it hit tree and comes to rest, then velocity of the car, "v = 0 m\/s"

Change in kinetic energy, "\Delta K.E. = \frac{1}{2}mu^2-\frac{1}{2}mv^2 = \frac{1}{2}mu^2 = \frac{1}{2}\times 1720\times (10.3)^2 = 91.24\times 10^3 J"


Work done by the car is equal to change in kinetic energy,

"W = \Delta K.E. = 91.24 \times 10^3 J"


Force is find by the formula,

"W = Fd"

where d is the displacement (i.e. push)

"F = \frac{W}{d} = \frac{91.24 \times 10^3 }{0.5} = 182.5 \times 10^3 N"